OpenStudy (anonymous):

What are the next three terms of the sequence 1, -3, 9, -27,…? A. -54. 162, -486 B. -81, 243, -729 C. 81, -243, 729 D. 57, -162, 486

OpenStudy (anonymous):

I don't know this one sorry:(

OpenStudy (_michaeld22):

Seems like it's multiplying by (-3) each time. 1*(-3) = -3. -3*(-3) = 9. 9*(-3) = -27. -27*(-3) = ?

OpenStudy (marissalovescats):

Yes they are. The pattern is that each number is being multiplied by -3

OpenStudy (anonymous):

\(\Large 1, -3, 9, -27= (-3)^0, (-3)^1 , (-3)^2 , (-3)^3 \) So,the next three terms of the sequence: \(\Large (-3)^4,(-3)^5,(-3)^6=81,-243,729\) The correct answer is C